Subtract, answer in simplest form:
2/5 - y/y-2
Is it -3y-4/y-2?

Answer by fractalier(2101)

You can use the criss-cross method for this...A/B ± C/D = (AD ± BC) / BD...so we have
2/5 - y/y-2 =
[2(y-2) - 5y] / 5(y-2) =
(2y - 4 - 5y) / 5(y-2) =
(-3y - 4) / 5(y - 2)
I think you forgot the 5...